Shift report 20030304 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Mar 04 2003 - 00:19:02 EST

  • Next message: Apache: "Shift report 20030304 00:00-08:00"
    ------------------------------------------------------------
    Date: Tue Mar  4 00:19:02 2003, Shift: 16:00-24:00
    Supervisor: ykl, with: 
    Log book#14 pages 44 - 45
    
    Shift Summary:
    ------------------------------------------------------------
    1.Beam was on since 20:00 but the background was too high
    (TOFW/ZDC > 100). After several phone exchanges MCR gave up
    steering saying they might have some communication problem (FCC?)
    in the area.
    2. No achievements except pedestal, TPC-pedestal, and DVM runs,
    8192, 8195, 8193, respectively.
    3.D5 tripped. With setpoint of 1050 D5 measured current hovers
    around 1032.
    TOFII mainframe is in error state. Resetting TSIR terminal
    server would not allow to fix this.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Mar 04 2003 - 00:23:41 EST